votecity20:

A rectangular football pitch has its length equal to twice its width and a perimeter of 360 m. Find its length and width

sammixboo:

So here we are going to use the formula for the perimeter or a rectangle, and plug all of the information we have to solve. The formula is: P = 2(L x W) Let's plug in our information to get our new equation: 360 = 2(L x W) Now clearly we don't have enough information to solve at this moment. But we do know that the length is twice the width, so we can make that into an equation as well, and replace "L" with this new equation: L = 2w Now, we can replace "L" in our original equation with "2w" to get this: 360 = 2(2w + w) Do you know the next steps to solve this equation?

votecity20:

Yes Thank you

sammixboo:

Awesome! What did you get for the answer?

votecity20:

360/6=60 m and 120 m